### v1

@@ 4.3 Stage 3: Dissemination
“If it’s not accurate, especially if it’s done by someone [who’s] a fantastic artist, but they don’t have the science background to know that DNA is twisted the wrong way […] it looks really sexy, but it’s wrong—that’s a big concern.” Mary captures the confluence of skills that a BioMedVis expert applies to the development of a biomedical visualization. Specific design choices (how), even if aesthetic, are modulated by the purpose (why/for whom) and the data being visualized (what) (see Fig. 2). “With the scientific side of it, I think there is a bit of a difference between creating something that is meant to be truly artistic and something that is meant to be based on science,” asserts Margaret.
@@
Ultimately, the perspectives shared by participants genAI in the BioMedVis pipeline embody not just attitudes related to this technology but also core values and concerns about their work that transcend genAI. All participants share a commitment to scientific accuracy as one of the central concerns in BioMedVis, although developers express this commitment more in relation to the data while the designers express this as a commitment to the patient or other intended target audience. This latter perspective is reinforced under the theme relating to human-centred and empathetic design—here, we see the same values but articulated through different motivations. Values expressed by developers under this theme relate more to issues of trust and judgment around the data, while designers again primarily relate use and accountability to human as opposed to data concerns. Values related to creative expression, however, are similarly-rooted for both groups: designers and developers place high value on aesthetics, story, and novelty in their work, whether in developing novel shading algorithms, molecular dynamics animations, or didactic, painterly anatomy.
@@ 6 Discussion
In this section we synthesize findings from our respective pipeline and thematic analyses to explore the value landscape that we see underlying participants’ overarching attitudes towards genAI. This value landscape affords deeper understanding as to why some participants find genAI useful in certain instances while others do not. The uses for genAI in the BioMedVis workflow in some cases affirm the speculations posed by Schetinger et al. [62] while other instances are new—this is due largely to the new considerations introduced by biomedical data and our designer group of interviewees, whose standard output is more figurative than procedural. Our work continues the conversation within the community to understand possibilities and contexts of genAI use, alongside the tangible impacts it exerts on the everyday workflow for BioMedVis and visualization more broadly.
@@ 6.1 Exploring participants’ genAI value landscape
Characterizing participants in a value landscape for genAI use, we find that their attitudes and active use varied greatly. Participants localize as enthusiasts to skeptics (how participants think or feel about genAI) and adopters to avoiders (how participants are actually using genAI tools) according to several factors: perceived genAI utility in their personal workflows, personal values, e.g., openness to using this technology, negative or positive attitudes towards genAI in general, and whether they willingly choose to engage with the technology or rather feel pressured to. Other aspects such as work context and experience play a role, but differently than we expected: the biggest skeptics range from early-career industry professional to late-stage academic, and the strongest adopters are in middle- to later-stage careers in both academia and industry. Here, we outline and reason about the varied attitudes we observed amongst our participants, arrayed into five personas: Enthusiastic adopters, Curious adapters, Curious optimists, Cautious optimists, and Skeptical avoiders (as shown in Fig. 1).
@@
All participants described a need for control throughout the BioMedVis pipeline, but this control looks different depending on their attitude toward genAI. The skeptical avoiders reject genAI entirely, as they deem the risks to be too great. We also see a strong sense of ownership over their knowledge and cultivated skills that they do not feel the need to ‘augment.’ The cautious optimists still very much see their own direct, hands-on work as being central to their process, and feel encumbered by having to correct genAI outputs. They also questioned what would happen to their sense of creative fulfillment and joy were they to offload ‘lower-level’ tasks, like basic coding or rendering, to genAI. The curious optimists and curious adapters likewise value their craft, but are ready to adapt their workflows to use genAI for lower-level tasks or, interestingly, to generate content that is outside their core competency (e.g., sound, character design). The enthusiastic adopters describe their sense of control akin to that of a conductor, and so see genAI as a value add to their work.
@@
GenAI tools are developing rapidly to alter the visualization pipeline. To illustrate these changes and their effects on individuals in this field, we contrast study participants’ experiences with Schetinger et al. [62]’s speculations in their 2023 study. Fig. 2 captures our participants’ perspectives on current, possible, and unlikely genAI use in three areas: scientific accuracy for what is being visualized (blue); the task/need, or why, for developing a new method for, or producing, a visualization (red); and how creative choices drive the mechanics of visualization production (green). In each area, we discuss opportunities for collaboration with designers situated outside the core visualization field.
@@
Schetinger et al. [62] further predicted tasks related to visual design, including moodboarding, basic coding, rapid prototyping, and stylization or “beautification”, which were echoed in our study, although with a caveat: in an industry where unique style is key to winning client contracts, the bias toward novelty in BioMedVis generally limits the perceived utility of genAI, especially for many designers.
Copyright concerns regarding AI-generated images were mixed in Schetinger et al. [62]’s study. By contrast, BioMedVis designers do see their output as artistic artifacts, and so have a stronger sense of how genAI and copyright risk impact their work. As such, they offered a clearer delineation for appropriate use of AI-generated elements in the pipeline, i.e., earlier stages of development, and in later stages if minor or if modified. Tooling that assists designers in labelling when and how genAI is used in the pipeline would help address this issue.
Contrasting speculated vs. current support for design with nuance. Between our two studies, design for accessibility is seen as a meaningful application of genAI and is an active area of research (e.g. [75]). In BioMedVis, this benefits patient education and public health outreach. The challenge remains in assembling multiple elements (text, visuals, interactivity) into comprehensive pieces that communicate information effectively—a task for which genAI, according to our participants’ experience, is not ready.
@@ 6.3 A call to action for science communication
Thoughtful, nuanced design serves as an essential bridge to how our target audiences receive scientific content. In contemplating how we collectively reason about genAI, we frame human partnerships with genAI tools as an exercise in crafting, rather than finding or revealing, data [72]. This process of data crafting includes numerous touch points where we make creative choices and use contextual knowledge to produce a result that reflects our understanding of reality. Using genAI to facilitate the crafting of data and visuals eliminates many of these reflective opportunities for human choice in favour of machine decision, resulting in a synthesized product made of code and pixels that cannot be fully rationalized. For BioMedVis, this is problematic since spatial relationships and entangled variables require fine-grained rationalization be considered trustworthy, issues which can extend into the broader space of science communication when genAI tools are wielded to craft decisions rather than through careful choices.
@@
Provenance and validation tools for genAI outputs, including scientific and stylistic elements, are necessary to improve BioMedVis experts’ adoption of genAI. Close collaboration and prototyping with designers will be essential to understand the numerous touch points where these validation tools will be most useful in crafting data, and managing sources [32] in BioMedVis. There are other opportunities for visualization research to support genAI-assisted workflows, e.g., creating interfaces to semi-automate visual metaphor design and visual inspiration for moderately-complex anatomical structures. Data interpolation for medical animations (e.g., similar to [52]), or interpolation of 3D spatial data for more common structures, are also possible applications. Visualization research has drawn inspiration from illustrative visualization in years past, and we see this continuing through, e.g., style transfers for rendering anatomy, or in narrative visualization [66, 46] to translate visual styles according to different audiences for data-driven medical stories. Importantly, these must be steered by a human on an ethically-trained model.
@@
As genAI technologies integrate more seamlessly into creative and professional workflows, we find ourselves operating within an expanding grey area concerning what constitutes ethical use. How should we properly credit genAI when it contributes to our work? Conversely, is there a need to explicitly state when genAI has not been used? These questions are not just theoretical; they have real implications for how work is evaluated, both in terms of value and authenticity. For both developers and designers, this shift invites a reconsideration of how their work is perceived and valued. If part of the labour is shared with an genAI tool, how might our effort, originality, and worth of that output be assessed? Critical reflection of the ethical and legal boundaries within which we operate is essential as we move forward.
@@ 8 Conclusion
BioMedVis lies at the tension of science and figurative art, drawing from diverse disciplines to create visuals for health and medicine. As the latest disruptive technological innovation in the history of computing, genAI is poised to both support and hamstring this field. GenAI use is mediated by the BioMedVis field’s core commitment to scientific accuracy and visualizing the human body and its mechanisms with empathy. Concerns such as the way one and their work is perceived, the nature of work itself in this hyper-specialized field, and bottom-line financials, drive individual attitudes and (dis)use of genAI tools. Primarily used as an auxiliary workflow tool now, we find that, e.g., career stage, personal philosophies towards technology, and a deep sense of responsibility shape whether genAI is seen to craft decisions or facilitate careful human choices in service of scientific storytelling. In comparing current genAI use to earlier speculations, we continue the conversation around genAI as an assistive or disruptive agent for visualization. We further assert the need for continuous human intervention in a pipeline with potential for high-stakes consequences, should a sexy visual misrepresent essential health information. Opportunities for visualization research to support these aims and enhance genAI-assisted workflows can enhance broader adoption by the design community in particular. However, equally important to BioMedVis work are the relational, human aspects which should not be lost in our efforts toward accurate, trustworthy, and engaging scientific visuals.

### v2

@@ 4.3 Stage 3: Dissemination
“If it’s not accurate, especially if it’s done by someone [who’s] a fantastic artist, but they don’t have the science background to know that DNA is twisted the wrong way […] it looks really sexy, but it’s wrong—that’s a big concern.” Mary captures the confluence of skills that a BioMedVis expert applies to the development of a biomedical visualization. Specific design choices (how), even if aesthetic, are modulated by the purpose (why/for whom) and the data being visualized (what) (see Fig.˜2). “With the scientific side of it, I think there is a bit of a difference between creating something that is meant to be truly artistic and something that is meant to be based on science,” asserts Margaret.
@@
Ultimately, the perspectives shared by participants about genAI in the BioMedVis pipeline embody not just attitudes related to this technology but also core values and concerns about their work that transcend genAI. All participants share a commitment to scientific accuracy as one of the central concerns in BioMedVis, although developers express this commitment more in relation to the data while the designers express this as a commitment to the patient or other intended target audience. This latter perspective is reinforced under the theme relating to human-centred and empathetic design—here, we see the same values but articulated through different motivations. Values expressed by developers under this theme relate more to issues of trust and judgment around the data, while designers again primarily relate use and accountability to human as opposed to data concerns. Values related to creative expression, however, are similarly-rooted for both groups: designers and developers place high value on aesthetics, story, and novelty in their work, whether in developing novel shading algorithms, molecular dynamics animations, or didactic, painterly anatomy.
@@ 6 Discussion
In this section, we synthesize findings from our respective pipeline and thematic analyses to explore the value landscape that we see underlying participants’ overarching attitudes towards genAI. This value landscape affords deeper understanding as to why some participants find genAI useful in certain instances while others do not. The uses of genAI in the BioMedVis workflow, in some cases, affirm the speculations posed by Schetinger et al. [62], while other instances are new—this is due largely to the new considerations introduced by biomedical data and our designer group of interviewees, whose standard output is more figurative than procedural. Our work continues the conversation within the community to understand possibilities and contexts of genAI use, alongside the tangible impacts it exerts on the everyday workflow for BioMedVis and visualization more broadly.
@@ 6.1 Exploring participants’ genAI value landscape
Characterizing participants in a value landscape for genAI use, we find that their attitudes and active use varied greatly. Participants localize as enthusiasts to skeptics (how participants think or feel about genAI) and adopters to avoiders (how participants are actually using genAI tools) according to several factors: perceived genAI utility in their personal workflows, personal values, e.g., openness to using this technology, negative or positive attitudes towards genAI in general, and whether they willingly choose to engage with the technology or rather feel pressured to. Other aspects such as work context and experience play a role, but differently than we expected: the biggest skeptics range from early-career industry professional to late-stage academic, and the strongest adopters are in middle- to later-stage careers in both academia and industry. Here, we outline and reason about the varied attitudes we observed amongst our participants, arrayed into five personas: Enthusiastic adopters, Curious adapters, Curious optimists, Cautious optimists, and Skeptical avoiders (as shown in Fig.˜1).
@@
All participants described a need for control throughout the BioMedVis pipeline, but this control looks different depending on their attitude toward genAI. The skeptical avoiders reject genAI entirely, as they deem the risks to be too great. We also see a strong sense of ownership over their knowledge and cultivated skills that they do not feel the need to ‘augment.’ The cautious optimists still very much see their own direct, hands-on work as being central to their process, and feel encumbered by having to correct genAI outputs. They also question what would happen to their sense of creative fulfillment and joy were they to offload ‘lower-level’ tasks, like basic coding or rendering, to genAI. The curious optimists and curious adapters likewise value their craft, but are ready to adapt their workflows to use genAI for lower-level tasks or, interestingly, to generate content that is outside their core competency (e.g., sound, character design). The enthusiastic adopters describe their sense of control akin to that of a conductor, and so see genAI as a value add to their work.
@@
GenAI tools are developing rapidly to alter the visualization pipeline. To illustrate these changes and their effects on individuals in this field, we contrast study participants’ experiences with Schetinger et al. [62]’s speculations in their 2023 study. Fig.˜2 captures our participants’ perspectives on current, possible, and unlikely genAI use in three areas: scientific accuracy for what is being visualized (blue); the task/need, or why, for developing a new method for, or producing, a visualization (red); and how creative choices drive the mechanics of visualization production (green). In each area, we discuss opportunities for collaboration with designers situated outside the core visualization field.
@@
Schetinger et al. [62] further predicted tasks related to visual design, including moodboarding, basic coding, rapid prototyping, and stylization or “beautification,” which were echoed in our study, although with a caveat: in an industry where unique style is key to winning client contracts, the bias toward novelty in BioMedVis generally limits the perceived utility of genAI, especially for many designers.
Copyright concerns regarding AI-generated images were mixed in Schetinger et al. [62]’s study. By contrast, BioMedVis designers do see their output as artistic artifacts, and so have a stronger sense of how genAI and copyright risk impact their work. As such, they offered a clearer delineation for appropriate use of AI-generated elements in the pipeline, i.e., in earlier stages of development, and in later stages only if minor or modified. Tooling that assists designers in labelling when and how genAI is used in the pipeline would help address this issue.
Contrasting speculated vs. current support for design with nuance. Between our two studies, design for accessibility is seen as a meaningful application of genAI and is an active area of research (e.g., [75]). In BioMedVis, this benefits patient education and public health outreach. The challenge remains in assembling multiple elements (text, visuals, interactivity) into comprehensive pieces that communicate information effectively—a task for which genAI, according to our participants’ experience, is not ready.
@@ 6.3 A call to action for science communication
Thoughtful, nuanced design serves as an essential bridge to how our target audiences receive scientific content. In contemplating how we collectively reason about genAI, we frame human partnerships with genAI tools as an exercise in crafting, rather than finding or revealing, data [72]. This process of data crafting includes numerous touch points where we make creative choices and use contextual knowledge to produce a result that reflects our understanding of reality. Using genAI to facilitate the crafting of data and visuals eliminates many of these reflective opportunities for human choice in favour of machine decision, resulting in a synthesized product made of code and pixels that cannot be fully rationalized. For BioMedVis, this is problematic since spatial relationships and entangled variables require fine-grained rationalization to be considered trustworthy, issues which can extend into the broader space of science communication when genAI tools are wielded to craft decisions rather than through careful choices.
@@
Provenance and validation tools for genAI outputs, including scientific and stylistic elements, are necessary to improve BioMedVis experts’ adoption of genAI. Close collaboration and prototyping with designers will be essential to understand the numerous touch points where these validation tools will be most useful in crafting data, and managing sources [32] in BioMedVis. There are other opportunities for visualization research to support genAI-assisted workflows, e.g., creating interfaces to semi-automate visual metaphor design and visual inspiration for moderately-complex anatomical structures. Data interpolation for medical animations (e.g., similar to [52]), or interpolation of 3D spatial data for more common structures, are also possible applications. Visualization research has drawn inspiration from illustrative visualization in years past, and we see this continuing through, e.g., style transfers for rendering anatomy, or in narrative visualization [66, 46] to translate visual styles according to different audiences for data-driven medical stories. Importantly, these must be steered by a human overseeing an ethically-trained model.
@@
As genAI technologies integrate more seamlessly into creative and professional workflows, we find ourselves operating within an expanding grey area concerning what constitutes ethical use. How should we properly credit genAI when it contributes to our work? Conversely, is there a need to explicitly state when genAI has not been used? These questions are not just theoretical; they have real implications for how work is evaluated, both in terms of value and authenticity. For both developers and designers, this shift invites a reconsideration of how their work is perceived and valued. If part of the labour is shared with a genAI tool, how might our effort, originality, and worth of that output be assessed? Critical reflection of the ethical and legal boundaries within which we operate is essential as we move forward.
@@ 8 Conclusion
BioMedVis lies at the tension of science and figurative art, drawing from diverse disciplines to create visuals for health and medicine. As the latest disruptive technological innovation in the history of computing, genAI is poised to both support and hamstring this field. GenAI use is mediated by the BioMedVis field’s core commitment to scientific accuracy and visualizing the human body and its mechanisms with empathy. Concerns such as the way one and their work is perceived, the nature of work itself in this hyper-specialized field, and bottom-line financials, drive individual attitudes and (dis)use of genAI tools. Primarily used as an auxiliary workflow tool now, we find that, e.g., career stage, personal philosophies towards technology, and a deep sense of responsibility shape whether genAI is seen to craft decisions or facilitate careful human choices in service of scientific storytelling. In comparing current genAI use to earlier speculations, we continue the conversation around genAI as an assistive or disruptive agent for visualization. We further assert the need for continuous human intervention in a pipeline with potential for high-stakes consequences, should a “sexy” visual misrepresent essential health information. Opportunities for visualization research to support these aims and enhance genAI-assisted workflows can enhance broader adoption by the design community in particular. However, equally important to BioMedVis work are the relational, human aspects, which should not be lost in our efforts toward accurate, trustworthy, and engaging scientific visuals.